OpenStudy (anonymous):

Arcsin(sin(2π)/3) = (2π)/3 True or false i dont see how it can be but ive been wrong before.

OpenStudy (acal21):

true impute in calculator

zepdrix (zepdrix):

Normally this is true, A composition of a function and it's inverse should give you the argument back. Example:\[\Large\rm f(f^{-1}(x))=x\] But when we're dealing with inverse trig functions, they have their range restricted.

zepdrix (zepdrix):

Arcsine is restricted to quadrants 1 and 4. So it will give us back the reference angle of 2pi/3 in the first quadrant.

OpenStudy (anonymous):

Reread it, the fraction is outside of the sine function

zepdrix (zepdrix):

Oh the 3 is outside? Oh interesting :o

OpenStudy (anonymous):

Could be a typo on their side, but easier to argue later than "oh well i assumed you made a typo"

zepdrix (zepdrix):

\[\Large\rm \arcsin\left[\frac{\color{orangered}{\sin(2\pi)}}{3}\right]=\arcsin\left[\frac{\color{orangered}{0}}{3}\right]=\arcsin\left[0\right]\]Arcsine of 0 is simply..... 0, yes? :)

OpenStudy (anonymous):

Yep, so it has to be false

zepdrix (zepdrix):

Even if it was a typo, it's false in both cases :) So you should be ok either way.

OpenStudy (anonymous):

thanks!:)